# Gram-like matrix preserving extensions of noncommutative polynomials to sum of Hermitian Squares

### v1

@@ 3. Extending a noncommutative polynomial to a SOHS
In this section, we address the following question, which captures the main theme of this article.
Question 1.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial.
  1. (a)
    What is the procedure to get a SOHS extension $\tilde{f}$ of $f$?
  2. (b)
    If $f$ has some SOHS parts, can it be extended to a SOHS $\tilde{f}$ keeping the gram-like matrices of the the SOHS parts of $f$ intact?
We now prepare the ground for defining Gram-like matrices, which will help us answering Question 1.
@@ 3.1. Gram-like matrices for symmetric and SOHS noncommutative polynomials
Let $G$ be a $k\times k$ positive semidefinite matrix and $W_{k}$ be a column vector of size $k$ with arbitrary monomial entries (not necessarily in lexicographical order). Then
@@
    Moreover, if no row (or column respectively) of $G$ is $\begin{pmatrix}0&0&\cdots&0\end{pmatrix}$ (or $\begin{pmatrix}0&0&\cdots&0\end{pmatrix}^{t}$ respectively), then degree of the polynomial $W_{k}^{\ast}\;G\;W_{k}$ is twice the degree of any of the highest degree monomials in $W_{k}$.
@@
Let $\zeta_{i}$, $1\leq i\leq k$, be any $k$ many monomials taken from $\mathbb{R}\langle\underline{X}\rangle$. Let
$W_{k}=\begin{pmatrix}\zeta_{1}&\zeta_{2}&\cdots&\zeta_{k}\end{pmatrix}^{t}\;.$
There exist some $i_{0}$ between $1$ and $k$ such that for all $1\leq i\leq k$,
${\rm deg}(\zeta_{i})\leq{\rm deg}(\zeta_{i_{0}})=d\;(\text{say}).$
let us now form $W_{s(d,n)}^{\text{unordered}}$ by adding remaining $s(d,n)-k$ many monomials $\eta_{j}$s, $1\leq j\leq s(d,n)-k$ from $\mathbb{R}\langle\underline{X}\rangle$ of degree at most $d$ with $W_{k}$, i.e,
$W_{s(d,n)}^{\text{unordered}}=\begin{pmatrix}\zeta_{1}&\cdots&\zeta_{k}&\eta_{1}&\cdots&\eta_{s(d,n)-k}\end{pmatrix}^{t}\;.$
Let $G^{\prime}$ be the $s(d,n)\times s(d,n)$ matrix given as follows:
@@
We then have,
$W_{k}^{\ast}\;G\;W_{k}={W_{s(d,n)}^{\text{unordered}}}^{\ast}\;G^{\prime}\;W_{s(d,n)}^{\text{unordered}}.$ (6)
Clearly, there exists a permutation $\sigma$ in the symmetric group $S_{s(d,n)}$ on $s(d,n)$ many symbols such that entries of the new column vector $W_{s(d,n)}$ (say), obtained by applying $\sigma$ to the rows of $W_{s(d,n)}^{\text{unordered}}$, is in lexicographic order. By $G^{\prime\prime}$ we denote the matrix obtained from $G^{\prime}$ by permuting its rows and columns by $\sigma$. Then we have,
${W_{s(d,n)}^{\text{unordered}}}^{\ast}\;G^{\prime}\;W_{s(d,n)}^{\text{unordered}}=W_{s(d,n)}^{\ast}\;G^{\prime\prime}\;W_{s(d,n)}.$ (7)
Clearly, the matrix $G^{\prime}$ is a positive semidefinite matrix as $G$ is so and by Lemma 2.7. Therefore, $G^{\prime\prime}$ is also a positive semidefinite matrix by Lemma 2.6 and the fact that $\sigma$ is composition of transpositions. Hence the first part of the assertion follows from Theorem 2.12, (6) and (7).
Let $\zeta_{i_{0}}$ be one of the monomials in $W_{k}$ of maximum degree. Then, by hypothesis, the $i$-th column and $i$-th row is nonzero. Therefore, by Lemma 2.8, the $(i_{0},i_{0})$-th entry of $G$ is positive, say $a_{i_{0}}$. Then, $f$ contains the term $a_{i_{0}}\zeta^{\ast}\zeta$ of degree $2{\rm deg}(\zeta_{i_{0}})$. So, ${\rm deg}(f)\geq 2{\rm deg}(\zeta_{i_{0}})$. The inequality ${\rm deg}(f)\leq 2{\rm deg}(\zeta_{i_{0}})$ is obvious as $W_{d}$ contains no monomial of degree bigger than ${\rm deg}(\zeta_{i_{0}})$. Hence, the second part of the assertion follows. ∎
@@
Let $G$ be any $k\times k$ symmetric matrix and $W_{k}$ be any column vector of size $k$ with arbitrary monomial entries (not necessarily in lexicographical order). Then the polynomial $W_{k}^{\ast}\;G\;W_{k}$ is a symmetric polynomial.
@@
  3. (3)
    Recall that Remark 2.13 provides an example of the fact that writing a SOHS polynomial in the form $W_{k}^{\ast}\;G\;W_{k}$ doesn’t guarantee the nonnegativity of $G$. Here we provide a simpler example incorporating the idea of the example provided in $(2)$ of Remark 3.3. Clearly, $f=2X_{1}X_{2}^{2}X_{1}$ is a SOHS as $f=(\sqrt{2}X_{2}X_{1})^{\ast}(\sqrt{2}X_{2}X_{1})$. For $W_{2}=\begin{pmatrix}1&X_{1}X_{2}^{2}X_{1}\end{pmatrix}^{t}$ and $G=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, $f=W_{2}^{\ast}\;G\;W_{2}$. But clearly $G$ is not positive semidefinite.
@@
Let $G$ be a symmetric matrix with positive diagonal entries and $W_{k}$ be a column vector of monomials from $\mathbb{R}\langle\underline{X}\rangle$. Then the polynomial $f:=W_{k}^{\ast}\;G\;W_{k}$ is a symmetric polynomial of degree $2d$, $d$ being the degree of any of the highest degree monomial entries of $W_{k}$.
@@
    Let $G$ be a $k\times k$ positive semidefinite matrix with no row (or column) being zero and $W_{k}$ be a column vector of size $k$ with arbitrary monomial entries of degree at most $d$. Then the polynomial $f=W_{k}^{\ast}\;G\;W_{k}$ is a SOHS polynomial of degree $2d$ with a representation $\mathbf{R}_{f}$ given as follows :
@@
    for any $\begin{split}&\text{for\;any\;}K=\{i_{0},i_{1},\ldots,i_{\mid K\mid}\}\subset\{1,2,\ldots,k\}\text{\;with\;}i_{0}<i_{1}<\cdots<i_{\mid K\mid}\\ &\text{the\;matrix\;}\begin{pmatrix}a_{i_{1}}&b_{i_{1}i_{2}}&\cdots&b_{i_{1}i_{\mid K\mid}}\\ b_{i_{1}i_{2}}&a_{i_{2}}&\cdots&b_{i_{2}i_{\mid K\mid}}\\ \vdots&\vdots&\ddots&\vdots\\ b_{i_{1}i_{\mid K\mid}}&b_{i_{2}i_{\mid K\mid}}&\cdots&a_{i_{\mid K\mid}}\end{pmatrix}\text{is\; positive\;semidefinite\;.}\end{split}$ (9c)
@@
    Conversely, let $f$ be a SOHS polynomial of degree $2d$ with a representation $\mathbb{R}_{f}$, as in (8), satisfying the conditions as given in (9). Then there exists a $k\times k$ positive semidefinite matrix $G$ with no row (or column) being zero and a column vector $W_{k}$ of size $k$ with monomial entries of degree at most $d$ such that $f=W_{k}^{\ast}\;G\;W_{k}$.
@@
The condition as in (9b) ensures that no column or row are full zero, that is,
@@
by Proposition 3.4 and Lemma 2.8. Moreover, the condition as in (9c) ensures that the matrix whose $(i,j)$-th entries are $b_{ij}$ for all $i\neq j$, and $(i,i)$-th entries are $a_{i}$ for all $i$, is positive semidefinite by Theorem 2.5. Now it follows from Proposition 3.1. ∎
@@
    Let $G$ be a $k\times k$ symmetric matrix with positive diagonal entries and $W_{k}$ be a column vector of size $k$ with arbitrary monomial entries of degree at most $d$. Then the polynomial $f=W_{k}^{\ast}\;G\;W_{k}$ is a symmetric polynomial of degree $2d$ with a representation $\mathbf{R}_{f}$ given as follows :
@@
    Conversely, let $f$ be a symmetric polynomial of degree $2d$ with a representation $\mathbf{R}_{f}$, as in (10), satisfying the conditions as given in (11). Then there exists a $k\times k$ symmetric matrix $G$ with positive diagonal entries and a column vector $W_{k}$ of size $k$ with monomial entries of degree at most $d$ such that $f=W_{k}^{\ast}\;G\;W_{k}$.
@@
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a symmetric polynomial of degree $2d$ in $n$ variables and $k$ be any positive integer satisfying $k\leq s(d,n)$. If there exist a $k\times k$ symmetric matrix $G$ with positive diagonal entries such that $f$ can be written in the form $W_{k}^{\ast}\;G\;W_{k}$ for some column vector $W_{k}$ of monomials of degree at most $d$, then $G$ is called a Gram-like matrix of $f$ (associated with $\mathbf{R}_{f}$, as in (10), along with conditions as in (11)).
@@
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a SOHS polynomial of degree $2d$ in $n$ variables and $k$ be any positive integer satisfying $k\leq s(d,n)$. If there exists a $k\times k$ positive semidefinite matrix $G$ such that $f$ can be written in the form $W_{k}^{\ast}\;G\;W_{k}$ for some column vector $W_{k}$ of monomials of degree at most $d$, then $G$ is called a positive semidefinite Gram-like matrix of $f$ (associated with $\mathbf{R}_{f}$, as in (8), along with conditions as in (9)).
@@
    The difference between a Gram matrix and a Gram-like matrix of a polynomial can be seen through the following example. A positive semidefinite Gram-like matrix of the polynomial $f=X_{1}^{2}+4X_{1}X_{2}+4X_{2}X_{1}+16X_{2}^{2}$, as in Example 2.14, is $G^{\prime\prime}=\begin{pmatrix}16&4\\ 4&1\end{pmatrix}$, whereas each of its positive semidefinite Gram matrix is of size $7$. One of its positive semidefinite Gram matrix $G_{f}$ is given as follows
    $G_{f}=\begin{pmatrix}0&0&0&0&0&0&0\\ 0&1&0&0&4&0&0\\ 0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0\\ 0&4&0&0&16&0&0\\ 0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0\end{pmatrix}\;.$
@@
    The motivation for working with a Gram-like matrix of a noncommutative polynomial $f$ instead of its Gram matrix is to avoid unnecessary monomials in the corresponding monomial vector, which is same as ignoring any full zero row (or column) from a Gram matrix. The transition from a Gram matrix to a Gram-like matrix can be interpreted through the lens of graph theory as well. To be specific, the sparsity pattern (cf. [24, §8.1, p. 329-330] for more details) of a Gram-like matrix is always connected, whereas it is not so for a Gram matrix in general.
3.2. Block Matrix Technique
In this subsection, we answer Question 1 in various setup using “block matrix technique” by exploiting the flexibility obtained through Theorem 3.5 and Theorem 3.6. In due course, we prevail this technique.
Let us begin with the most basic setup. Given any noncommutative SOHS $f$, it is easy to observe that $f+c$ is also a SOHS for any $c>0$. As a warm up, we again prove this incorporating the mentioned block matrix technique.

### v2

@@ 3. Extending a noncommutative polynomial to a SOHS
In this section, whenever we consider a noncommutative polynomial $f\in\mathbb{R}\langle\underline{X}\rangle$, where $\langle\underline{X}\rangle=\langle X_{1},X_{2},\dots,X_{n}\rangle$, we mean that all the variables $X_{i}$, $1\leq i\leq n$, are essential to express $f$. Moreover, we want to find a SOHS extension $\widetilde{f}$ of $f$ in $\mathbb{R}\langle\underline{X}\rangle$ only. Specifically, we now prepare the ground for defining Gram-like matrices, which will help us answer the Question 1 that captures the main theme of this article.
@@ 3.1. Gram-like matrices for symmetric and SOHS noncommutative polynomials
Let $G$ be a $k\times k$ positive semidefinite matrix and $W_{k}$ be a column vector of size $k$ with $k$ many distinct monomial entries (not necessarily in lexicographical order). Then
@@
    Moreover, if no row (and therefore column) of $G$ is zero, then degree of the polynomial $W_{k}^{\ast}\;G\;W_{k}$ is twice the degree of any of the highest degree monomials in $W_{k}$.
@@
  1. (1)
    Let $d$ be the highest degree among the degrees of the constituent monomials of $W_{k}$. Let $W_{s(d,n)}$ be the monomial vector formed by taking the remaining $s(d,n)-k$ many monomials of degree at most $d$ along with the monomials that are already present in $W_{k}$ and then rearranging them in lexicographic order. This rearrangement is nothing but a permutation between $\{1,2,\cdots,s(d,n)\}$ to itself and we denote it by $\sigma\in S_{s(d,n)}$, the symmetric group on $s(d,n)$ many symbols. Let $G^{\prime}$, a $s(d,n)\times s(d,n)$ matrix, be defined as follows:
@@
    Let ${G^{\prime}}^{\sigma}$, another $s(d,n)\times s(d,n)$ matrix, be obtained from $G^{\prime}$ by permuting its rows and columns by $\sigma$. Then ${G^{\prime}}^{\sigma}$ is positive semidefinite as $G$ is so and by Lemma 2.9 & Lemma 2.8. Moreover, we have the following equality:
    $W_{k}^{\ast}\;G\;W_{k}=W_{s(d,n)}^{\ast}\;{G^{\prime}}^{\sigma}\;W_{s(d,n)}\;.$
    Therefore, the assertion follows from Theorem 2.16.
  2. (2)
    Let $\zeta_{i_{0}}$ be one of the monomials in $W_{k}$ of maximum degree. Then, by hypothesis, the $i_{0}$-th column and $i_{0}$-th row is nonzero. Therefore, by Lemma 2.10, the $(i_{0},i_{0})$-th entry of $G$ is positive, say $a_{i_{0}}$. Then, $f$ contains the term $a_{i_{0}}\zeta_{i_{0}}^{\ast}\zeta_{i_{0}}$ of degree $2{\rm deg}(\zeta_{i_{0}})$. So, ${\rm deg}(f)\geq 2{\rm deg}(\zeta_{i_{0}})$. The inequality ${\rm deg}(f)\leq 2{\rm deg}(\zeta_{i_{0}})$ is obvious as $W_{k}$ contains no monomial of degree bigger than ${\rm deg}(\zeta_{i_{0}})$. Hence, the assertion follows.
@@
Let $G$ be any $k\times k$ symmetric matrix and $W_{k}$ be any column vector of size $k$ with $k$ many distinct monomial entries (not necessarily in lexicographical order). Then the polynomial $W_{k}^{\ast}\;G\;W_{k}$ is a symmetric polynomial.
@@
Let $G$ be a symmetric matrix with positive diagonal entries and $W_{k}$ be a column vector of distinct monomials from $\mathbb{R}\langle\underline{X}\rangle$. Then the polynomial $f:=W_{k}^{\ast}\;G\;W_{k}$ is a symmetric polynomial of degree $2d$, $d$ being the degree of any of the highest degree monomial entries of $W_{k}$.
@@
    Let $G$ be a $k\times k$ positive semidefinite matrix with no row (or column) being zero and $W_{k}$ be a column vector of size $k$ with $k$ many distinct monomial entries of degree at most $d$. Then the polynomial $f=W_{k}^{\ast}\;G\;W_{k}$ is a SOHS polynomial of degree $2d$ with a representation $\mathbf{R}_{f}$ given as follows :
@@
    for any $\begin{split}&\text{for\;any\;}K=\{i_{0},i_{1},\ldots,i_{\mid K\mid}\}\subset\{1,2,\ldots,k\}\text{\;with\;}i_{0}<i_{1}<\cdots<i_{\mid K\mid}\\ &\text{the\;matrix\;}\begin{pmatrix}a_{i_{1}}&b_{i_{1}i_{2}}&\cdots&b_{i_{1}i_{\mid K\mid}}\\ b_{i_{1}i_{2}}&a_{i_{2}}&\cdots&b_{i_{2}i_{\mid K\mid}}\\ \vdots&\vdots&\ddots&\vdots\\ b_{i_{1}i_{\mid K\mid}}&b_{i_{2}i_{\mid K\mid}}&\cdots&a_{i_{\mid K\mid}}\end{pmatrix}\text{is\; positive\;semidefinite\;.}\end{split}$ (7d)
@@
    Conversely, let $f$ be a SOHS polynomial of degree $2d$ with a representation $\mathbf{R}_{f}$, as in (6), satisfying the conditions as given in (7). Then there exists a $k\times k$ positive semidefinite matrix $G$ with no row (or column) being zero and a column vector $W_{k}$ of size $k$ with distinct monomial entries of degree at most $d$ such that $f=W_{k}^{\ast}\;G\;W_{k}$.
@@
The condition as in (7c) ensures that no column or row are full zero, that is,
@@
by Proposition 3.4 and Lemma 2.10. Moreover, the condition as in (7d) ensures that the matrix whose $(i,j)$-th entries are $b_{ij}$ for all $i\neq j$, and $(i,i)$-th entries are $a_{i}$ for all $i$, is positive semidefinite by Lemma 2.7. Now it follows from Proposition 3.1. ∎
@@
    Let $G$ be a $k\times k$ symmetric matrix with positive diagonal entries and $W_{k}$ be a column vector of size $k$ with distinct monomial entries of degree at most $d$. Then the polynomial $f=W_{k}^{\ast}\;G\;W_{k}$ is a symmetric polynomial of degree $2d$ with a representation $\mathbf{R}_{f}$ given as follows :
@@
    $a_{i},b_{ij}\in\mathbb{R},\zeta_{i}\in\mathbb{R}\langle\underline{X}\rangle,\;\zeta_{i}\neq\zeta_{j},\;{\rm deg}(\zeta_{i})\leq d\text{\;for\;all\;}1\leq i<j\leq k\;,$ (9a)
@@
    Conversely, let $f$ be a symmetric polynomial of degree $2d$ with a representation $\mathbf{R}_{f}$, as in (8), satisfying the conditions as given in (9). Then there exists a $k\times k$ symmetric matrix $G$ with positive diagonal entries and a column vector $W_{k}$ of size $k$ with distinct monomial entries of degree at most $d$ such that $f=W_{k}^{\ast}\;G\;W_{k}$.
@@
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a symmetric polynomial of degree $2d$ in $n$ variables and $k$ be any positive integer satisfying $k\leq s(d,n)$. If there exist a $k\times k$ symmetric matrix $G$ with positive diagonal entries such that $f$ can be written in the form $W_{k}^{\ast}\;G\;W_{k}$ for some column vector $W_{k}$ of monomials of degree at most $d$, then $G$ is called a Gram-like matrix of $f$ (or the Gram-like matrix associated with $\mathbf{R}_{f}$, as in (8), along with conditions as in (9)).
@@
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a SOHS polynomial of degree $2d$ in $n$ variables and $k$ be any positive integer satisfying $k\leq s(d,n)$. If there exists a $k\times k$ positive semidefinite matrix $G$ such that $f$ can be written in the form $W_{k}^{\ast}\;G\;W_{k}$ for some column vector $W_{k}$ of monomials of degree at most $d$, then $G$ is called a positive semidefinite Gram-like matrix of $f$ (or the positive semidefinite Gram-like matrix associated with $\mathbf{R}_{f}$, as in (6), along with conditions as in (7)).
@@
    The difference between a Gram matrix and a Gram-like matrix of a polynomial can be seen through the following example. A positive semidefinite Gram-like matrix of the polynomial $f=X_{1}^{2}+4X_{1}X_{2}+4X_{2}X_{1}+16X_{2}^{2}$, as in Example 2.18, is $G^{\prime\prime}=\begin{pmatrix}16&4\\ 4&1\end{pmatrix}$, whereas each of its positive semidefinite Gram matrix is of size $7$.
@@
    The motivation for working with a Gram-like matrix of a noncommutative polynomial $f$ instead of its Gram matrix is to avoid unnecessary monomials in the corresponding monomial vector, which is same as ignoring any full zero row (or column) from a Gram matrix. The transition from a Gram matrix to a Gram-like matrix can be interpreted through the lens of graph theory as well. To be specific, the sparsity pattern (cf. [25, §8.1, p. 329-330] for more details) of a Gram-like matrix is always connected, whereas it is not so for a Gram matrix in general.
3.2. Gram-like matrix preserving SOHS extensions
In this subsection, we answer Question 1 completely by exploiting the flexibility obtained through Theorem 3.5 and Theorem 3.6. To do so, let us formally define the meaning of SOHS extension of a noncommutative polynomial.
Definition 3.10.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial of the form $f=\sum_{j=1}^{m}a_{j}\delta_{j}$, where $\delta_{j}$ are monomials in $\mathbb{R}\langle\underline{X}\rangle$, $a_{j}\in\mathbb{R}-\{0\}$ for all $j$. A SOHS polynomial $\widetilde{f}\in\mathbb{R}\langle\underline{X}\rangle$ is called a nontrivial SOHS extension of $f$ if $\widetilde{f}-f=\sum_{i=1}^{n}b_{i}\alpha_{i}$, where $\alpha_{i}$ are monomials in $\mathbb{R}\langle\underline{X}\rangle$, $\alpha_{i}\neq\delta_{j}$, $b_{i}\in\mathbb{R}-\{0\}$ for all $i,j$.
From now on, by a SOHS extension we mean a nontrivial SOHS extension only, as defined above. We now define a special type of SOHS extension, namely, a Gram-like matrix preserving extension.
Definition 3.11.
A SOHS extension $\widetilde{f}\in\mathbb{R}\langle\underline{X}\rangle$ of a noncommutative polynomial $f\in\mathbb{R}\langle\underline{X}\rangle$ is said to be a Gram-like matrix preserving extension of $f$ w.r.t $h$ if
  • there exists a $l\times l$ positive semidefinite Gram-like matrix $G_{\widetilde{f}}$ of $\widetilde{f}$ that keeps a $k\times k$ positive semidefinite Gram-like matrix $G_{h}$ of a SOHS part $h$ of $f$ as a principal submatrix, for some positive integers $l>k$, and
  • there exists a monomial vector $W_{l}$ associated to $G_{\widetilde{f}}$ and a monomial vector $W_{k}$ associated to $G_{h}$ such that $W_{l}$ contains all the entries of $W_{k}$.
Remark 3.12.
From now on, we simply say that $G_{\widetilde{f}}$ is a positive semidefinite Gram-like matrix of a Gram-like matrix preserving extension $\widetilde{f}$ of $f$ to mean that such $G_{\widetilde{f}}$ satisfies both the conditions as in Definition 3.11.
Example 3.13.
  1. (1)
[…]
  2. (2)
    By definition, it is obvious that any Gram-like matrix preserving extension w.r.t some SOHS part of a given noncommutative polynomial is a SOHS extension. However, the converse is not true. Consider the polynomial $f=X^{2}+X+5$ with $h=X^{2}+5$. Then $\widetilde{f}=X^{2}+X+5+X^{4}=f+X^{4}$ is a SOHS extension of $f$. But, as explained above, $f$ doesn’t admit a Gram-like matrix preserving extension w.r.t $h$.
The following lemma talks about the general form of a positive semidefinite Gram-matrix of a Gram-like matrix preserving extension.
Lemma 3.14.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial and $h$ a SOHS part of $f$. Let $\widetilde{f}$ be a Gram-like matrix preserving extension of $f$ w.r.t $h$ and $G_{h}$ and $G_{\widetilde{f}}$ be as in Definition 3.11. Then, up to permutation, $G_{\widetilde{f}}$ must be of the form
$G_{\widetilde{f}}=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}G_{h}\end{matrix}&A\\ \hline\cr A^{t}&\begin{matrix}B\end{matrix}\end{array}\right)\;,$
where $G_{\widetilde{f}}/A(:=A^{t}B^{-}A)$ is positive semidefinite and $\mathcal{C}(A)\subseteq\mathcal{C}(G_{h})$.
Proof.
Follows from Definition 3.11, Lemma 2.8 and Lemma 2.12. ∎
We now look for some necessary conditions and sufficient conditions for the existence of a Gram-like matrix preserving extension of a given noncommutative polynomial $f$, in terms of the monomials of $f$. We set the stage for that.
The following lemma talks about how a given monomial can be decomposed through two smaller sized monomials.
Lemma 3.15.
Any monomial $W\in\langle\underline{X}\rangle$ can be decomposed as a product of two monomials in ${\rm deg}(W)+1$ many ways.
Proof.
Any monomial $W\in\langle\underline{X}\rangle$ can be decomposed as following:
$W=\big{(}{\rm rc}(W^{*},{\rm deg}(W)-n)\big{)}^{*}{\rm rc}(W,n),$ (10)
for any integer $n$ satisfying $0\leq n\leq{\rm deg}(W)$ and therefore there are ${\rm deg}(W)+1$ many such ways. ∎
Definition 3.16.
Given a monomial $W\in\langle\underline{X}\rangle$, any decomposition of $W$, as in (10), is called a rc deomposition of $W$ (Right Chip decomposition of $W$).
The following lemma provides a way to produce a positive semidefinite matrix out of a partially specified symmetric matrix of a particular type. This is very crucial in proving the existence of a Gram-like matrix preserving extension.
Lemma 3.17.
Let
$G=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}A\end{matrix}&B\\ \hline\cr B^{t}&\begin{matrix}C\end{matrix}\end{array}\right)\;$
be a partially specified symmetric matrix, where $A$ and $C$ are square matrices of size $m$ and $n$ respectively and $A$ is a positive semidefinite matrix. Let the diagonal elements of $C$ be the only unspecified entries of $G$. Then if the column space of $B$ sits inside the column space of $A$, that is, $\mathcal{C}(B)\subseteq\mathcal{C}(A)$, then it is possible to fill the diagonal elements of $C$ with some positive real numbers so that the matrix $G$ becomes positive semidefinite.
Proof.
By Lemma 2.12, it is enough to choose diagonals $c_{ii}$ of $C$ such that
  • $c_{ii}>0$ for all $1\leq i\leq n$,
  • The matrix $G/A:=C-B^{t}A^{-}B$ is positive semidefinite.
We choose $c_{ii}>0$ satisfying $c_{ii}\geq\sum_{j\neq i}|(C-B^{t}A^{-}B)_{ij}|$, for all $i$. Clearly, such choices make the matrix $G/A:=C-B^{t}A^{-}B$ diagonally dominant, and hence positive semidefinite. Hence the assertion follows. ∎
Remark 3.18.
In Lemma 3.17, if the matrix $A$ is positive definite, then condition $\mathcal{C}(B)\subseteq\mathcal{C}(B)$ is automatically satisfied.
We are now ready to provide some necessary and sufficient conditions for the existence of a Gram-like matrix preserving extension.